OpenStudy (anonymous):

hey guys can sum1 please help me anwser this question thanks..calculate the y intercept (b) of a line that is perpendicular to the line y=8x+3 and that gose threw the point (-1,-3)

OpenStudy (anonymous):

Gradient of perpendicular line is -1/8. So y=-x/8+c, so -3=1/8+c, so c=-25/8

OpenStudy (anonymous):

i appreciate the reply so my anwser would be c=25/8??

OpenStudy (anonymous):

-25/8

OpenStudy (anonymous):

sry -25/8?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

u r the best i really hope this is correct

OpenStudy (anonymous):

do you kind helping me with another 1 or 2?

OpenStudy (anonymous):

Yeah, ok then.

OpenStudy (anonymous):

calculate the slope of the line that these two points lie on (-5,0) (1,-5)

OpenStudy (anonymous):

how would -25/8 be graphed though?

OpenStudy (anonymous):

\[m=\frac{\Delta y}{\Delta x}=\frac{y _{2}-y _{1}}{x _{2}-x _{1}}=\frac{-5}{6}\]

OpenStudy (anonymous):

What do you mean how would it be graphed?

OpenStudy (anonymous):

slope = (y2-y1)/(x1 - x2) = (-5-0)/[1-(-5)] = -5/6

OpenStudy (anonymous):

i dont even know im just really confused lol .. -25/8 is the y intersect?

OpenStudy (anonymous):

ty karkirat

OpenStudy (anonymous):

Yes which is -(3 & a eighth) if that helps.

OpenStudy (anonymous):

harkirat *

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

would u guys be able 2 help me with a matrix problem?

OpenStudy (anonymous):

if a = b= -3 -5 5 2 8 -2 -3 1 then what is the element in the first row first colum of the matrix X if; B+ (1/3)X=2A

OpenStudy (anonymous):

i just dont understand what x is coming from wat is x

OpenStudy (anonymous):

X is defined by the equation B_(1/3)X=2A.

OpenStudy (anonymous):

the dterminant if a is 46 d of b is 11

OpenStudy (anonymous):

so x would have to be 243

OpenStudy (anonymous):

but how do make x a matric in order to get the 1st row 1st colum element

OpenStudy (anonymous):

Ok, think about it like this: the top left element of 2A is -6, you need some number that when you divide it by 3 and add it to 5, you get -6. So the top left element of X is -33.

OpenStudy (anonymous):

would there be a way to solve that if i just havbe the determinant for matrix x?

OpenStudy (anonymous):

also i am submitting these anwsers online so in the anwser box for the previous 2 questions is i woulkd write -25/8 -5/6 would that be correct or it should be like (-5,6)

OpenStudy (anonymous):

-25/8 & -5/6

OpenStudy (anonymous):

You don't need to know the determinant of X because you know A & B.

OpenStudy (anonymous):

hey buddy r u still there?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

sry but do u mind helping me out again wat just happened to me is unbelievable i took a test online took me 4 hrs to complete and it did not go threw and i have a completly new 1 with differnt question

OpenStudy (anonymous):

Calculate the y-intercept (b) of a line that is perpendicular to the line y=2x+2 and that goes through the point (-3,4).

OpenStudy (anonymous):

Ok, how about I show you how to do these and you can work them out for yourself?

OpenStudy (anonymous):

can u prety please just help me out with like 3 i will do the rest i spent 4 hrs on the other test and have 2 solve many of these as well and have not even slept yet and have class

OpenStudy (anonymous):

Yeah, sure, I was just thinking it might be better for you if I show you the method rather than just doing each question :)

OpenStudy (anonymous):

please man i know thats the right thing but this is unbelievable i cant believe im still doing math at 8 am i have not slept

OpenStudy (anonymous):

Wow is it 8 o'clock there?! It's 1 in the afternoon here :P

OpenStudy (anonymous):

Ok, m=-1/2, so y=-x/2+c, so 4=3/2+c, so c=2.5.

OpenStudy (anonymous):

so the anwser for y intercept is 2.5?

OpenStudy (anonymous):

If -4 -8 A = 6 6 and 5 2 B = -3 1 then what is the element in the first row and first column of the matrix X if: B + (1/3)X = 2A

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

-54

OpenStudy (anonymous):

thanks now that matrix question plz

OpenStudy (anonymous):

No wait, that's not right. It's -36.

OpenStudy (anonymous):

take ur time buddy lol

OpenStudy (anonymous):

Sorry, -39, definitely -39.

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

By using Gauss Jordan Elimination find what z is equal to from the following set of three equations: x - y + 3z = 14 x + y + z = 6 -2x - y +z = 0

OpenStudy (anonymous):

Isn't Gauss-Jordan Elimination to do with matrices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats wat we have been doing no? lol

OpenStudy (anonymous):

I don't know what that has to do with it but z=4.

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

im done do u mind if i just give my anwsers so u can see if they r correct

OpenStudy (anonymous):

No problem, go for it.

OpenStudy (anonymous):

Find what y is equal to in the following system of equations by whichever method you prefer. 1x+4y=5 -3x+8y=1 Answer .8

OpenStudy (anonymous):

If 5 1 A = 8 -5 and 5 2 B = -3 -1 then what is the element in the second row and first column of the matrix that results from mulitplying matrix A times matrix B ? Answer 55

OpenStudy (anonymous):

If -4 -8 A = 6 6 and 5 2 B = -3 1 then what is the element in the first row and first column of the matrix X if: B + (1/3)X = 2A -39

OpenStudy (anonymous):

I make the first one -3.

OpenStudy (anonymous):

Find what y is equal to in the following system of equations by whichever method you prefer. 7x+-4y=4 3x+-3y=4 1.777

OpenStudy (anonymous):

Find the distance between the following two points: (-6,5) and (5,-3) 13.601

OpenStudy (anonymous):

If 2 -4 A = -3 -6 and 5 2 B = -3 1 then what is the element in the second row and second column of the matrix that results from subtracting matrix A from matrix B ? -7

OpenStudy (anonymous):

and thats itthanks man let me know wat u think

OpenStudy (anonymous):

In the system of equations i make it -16/9=-1.7777...

OpenStudy (anonymous):

Distance between points is correct.

OpenStudy (anonymous):

The matrix subtraction is 7.

OpenStudy (anonymous):

y r we gettin differnt signs for everything

OpenStudy (anonymous):

I don't know but that matrix one is definitely 7: 1-(-6)=7

OpenStudy (anonymous):

im using an online calculator it says -7cause its -6-1 not 1--6

OpenStudy (anonymous):

Subtracting A from B is B-A not A-B.